OpenStudy (anonymous):

write a quadratic equation having the given numbers as solutions. -10/3,2

OpenStudy (amistre64):

given numbers as solutions to what?

OpenStudy (caozeyuan):

to the qe, I think

OpenStudy (amistre64):

the term solution is ambiguous are those roots? or something else

OpenStudy (caozeyuan):

of course they are roots

OpenStudy (amistre64):

f(x) = -10/3 and f(x) = 2 are also defined by the term 'solution'

OpenStudy (anonymous):

i got 3x^2+3x-20=0 but im not sure if its right

OpenStudy (amistre64):

well, to dbl chk, insert the solutions to see of its true

OpenStudy (amistre64):

otherwise: a product of the roots would turn out: (-10/3 - x) (2 - x) (-10 - 3x) (2 - x) 3x^2 +20 -6x +10x

OpenStudy (amistre64):

id try 4x instead of 3x

OpenStudy (amistre64):

-20 ... my fingers hate me

OpenStudy (anonymous):

its 3x^2+4x-20=0 i had to redo it thanks for the help though

OpenStudy (amistre64):

youre welcome